OpenStudy (anonymous):

The equation of an ellipse is shown (I'll post it in a minute). What is the length of it's major axis?

OpenStudy (tkhunny):

Eyeball problem. 30 Just recognize the form. 225 = a^2 ==> a = 15 ==> 2*a = 30

OpenStudy (anonymous):

\[\frac{ (x-60)^2 }{ 256 } + \frac{ (y-34)^2 }{ 196 } = 1\]

OpenStudy (epoweritheta):

what is root of 256 ??

OpenStudy (epoweritheta):

16 is the length of major axis

OpenStudy (tkhunny):

?? \(\sqrt{256} = 16 = a\) \(Length\;of\;Major\;Axis = 2a = 2(16) = 32\) Again, this is an eyeball problem. Just look at it and write down the answer.